chandra is tossing a softball into the air with an underhand motion. the distance of the ball above her hand at any time is given by the function h(t)=32t-16t^2 for 0 < or equal to t < or equal to 2. where h (t) is the height and t is the time (in seconds). find the times at which the ball is in her hand and the maximum height of the ball.
Found 2 solutions by ewatrrr, ikleyn:
Answer by ewatrrr(24785)


Hi,
chandra is tossing a softball into the air with an underhand motion. 
the distance of the ball above her hand at any time is given by the function 

h(t)=32t-16t^2 for 0.  h(t) = 0 

    32t - 16t^2 = 0   (back down to the height of your hand)

    32t = 16t^2

            t = 2 sec   time it takes to get back down to the height of your hand

Wish You the Best in your Studies.


Answer by ikleyn(52879)

.

The height function is the parabola


     h(t) = -16t^ + 32t.


It has zeros, when  h(t) = 0,   or


    -16t^2 + 32t = 0.


Factor the expression


     -16t*(t-2) = 0.


The roots are  t= 0  and  t= 2.    (They are time moments, when the ball is at the ground level).


The maximum is reached at the time moment, which is exactly half-way between the roots: t = 1 second.    ANSWER


The height is maximum at this time


    h(t=1) = -16*1^2 + 32*1 = -16 + 32 = 16  feet.    ANSWER

Solved.

-----------------

In this site,  there is a bunch of lessons on a projectile thrown/shot/launched vertically up
    - Introductory lesson on a projectile thrown-shot-launched vertically up
    - Problem on a projectile moving vertically up and down
    - Problem on an arrow shot vertically upward
    - Problem on a ball thrown vertically up from the top of a tower
    - Problem on a toy rocket launched vertically up from a tall platform

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.
Read them attentively and learn how to solve this type of problems once and for all.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Projectiles launched/thrown and moving vertically up and dawn".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.